# GoMatching: A Simple Baseline for Video Text Spotting via Long and Short Term Matching

### v1

@@ 4.1 Datasets and Evaluation Metrics
ICDAR15-video Karatzas et al. (2015) is a word-level video text reading benchmark annotated with quadrilateral bounding boxes. It consists of a training set of 25 videos and a test set of 24 videos. ICDAR15-video focuses on wild scenarios, such as driving on the road, exploring shopping streets, walking in a supermarket, etc.
DSText Wu et al. (2023) is a newly proposed dataset, and focuses on dense and small text reading challenges in the video with various scenarios. This dataset provides 50 training videos and 50 test videos. Compared with the previous datasets, DSText mainly includes the following three new challenges: dense video texts, high-proportioned small texts, and various new scenarios, e.g., ‘Game’, ‘Sports’, etc. Similar to ICDAR15-video, DSText adopts word-level annotations, which are labeled with quadrilaterals.
ArTVideo is a novel test set established in this work to evaluate the performance of arbitrary-shaped video text. It contains 20 videos with 6,526 text instances including 4,632 straight text instances and 1,894 curved text instances (about 30% curved text instances). This dataset is annotated at the word level with quadrilaterals for straight text and polygons for curved text.
Evaluation Metrics. To evaluate performance, we adopt three evaluation metrics commonly used in ICDAR15-video competition and DSText competition, including MOTA (Multiple Object Tracking Accuracy) Bernardin and Stiefelhagen (2008), MOTP (Multiple Object Tracking Precision), and IDF1 Ristani et al. (2016).
@@ 4.2 Implementation Details
In all experiments, we only use a single NVIDIA GeForce RTX 3090 (24G) GPU to train and evaluate GoMatching. As for the image text spotter in GoMatching, we apply the officially released DeepSolo Ye et al. (2023a) trained on ICDAR15-image. During training GoMatching on ICDAR15-video and DSText, we only update the rescoring head and LST-Matcher, while keeping DeepSolo frozen.
Training Setting. For both ICDAR15-video and DSText, we follow EfficientDet Tan et al. (2020) to adopt the scale-and-crop augmentation strategy with the resolution of $280\times 1,$ 280 × 1 , 280. The batch size $T$ is 6. All frames in a batch are from the same video. Text instances with fusion scores higher than 0.3 are fed into the LST-Matcher during training. AdamW Loshchilov and Hutter (2017) is used as the optimizer. We adopt the warmup cosine annealing learning rate strategy with the initial learning rate being set to 0.00005. The loss weights $\lambda_{res}$ and $\lambda_{asso}$ are set to 1.0 and 0.5, respectively. For focal loss, $\alpha$ is 0.25 and $\gamma$ is 2.0 as in Carion et al. (2020); Ye et al. (2023a). The model is trained for 30k and 60k iterations on ICDAR15-video and DSText, respectively.
Inference Setting. Since there is no training set in ArTVideo, we directly use the model trained on ICDAR15-video to evaluate the generalization ability of GoMatching to arbitrary-shaped text. The shorter size of the input frame is resized to 1,440 on all datasets. We set the high-score instance threshold and NMS overlap threshold to 0.3 and 0.5 for ICDAR15-video, 0.5 and 0.3 for DSText, and 0.4 and 0.5 for ArTVideo, respectively. The association score threshold is set to 0.2.
@@ 4.3 Comparison with State-of-the-art Methods
Results on ICDAR15-video. To evaluate the effectiveness of GoMatching on oriented video text, we conduct a comparison with the state-of-the-art methods on ICDAR15-video presented in Table 1. As can be seen, GoMatching ranks first in all metrics on ICDAR15-video leaderboard. GoMatching presents excellent performance, compared to the second-best results, improving by 3.56% MOTA, 0.47% MOTP, and 1.75% IDF1, respectively. Furthermore, GoMatching outperforms current SOTA single-model method TransDETR by 9.56% MOTA, 3.64% MOTP, and 5.90% IDF1, respectively.
@@
GoMatching (ours) 70.52 78.25 78.70
Table 1: Results on ICDAR15-video. ‘†’ denotes that the results are collected from the official competition website. The best and second-best results are marked in bold and underlined, respectively.
@@
TransDETR Wu et al. (2022a)* N -22.63 79.73 26.43
GoMatching (ours) N 17.29 77.48 45.20
Table 2: Results on DSText. ‘†’ has the same meaning as in Table 1. ‘*’: we use the officially released model for evaluation. ‘M-ME’ indicates whether multi-model ensembling is used. ‘Y’ and ‘N’ stand for yes and no, respectively.
Results on DSText. We further conduct experiments on DSText to evaluate the effectiveness of GoMatching on dense and small video text scenarios. Results are presented in Table 2. It is worth noting that most of the previous methods on the DSText leaderboard utilized an ensemble of multiple models and large public datasets to enhance their performance Wu et al. (2023). For example, TencentOCR integrates the detection results of DBNet Liao et al. (2020b) and Cascade MaskRCNN Cai and Vasconcelos (2019) built with multiple backbone architectures, combines the Parseq model Bautista and Atienza (2022) for recognition, and further improves the end-to-end tracking with ByteTrack Zhang et al. (2022b). DA utilizes Mask R-CNN He et al. (2017) and DBNet to detect text, then uses BotSORT Aharon et al. (2022) to replace the tracker in VideoTextSCM Gao et al. (2021) and employs the Parseq model for recognition. As a single model with a frozen image text spotter, GoMatching also shows competitive performance compared to other ensembling methods on the leaderboard. GoMatching ranks second (17.29%) on MOTA and third (45.20%) on IDF1. Moreover, compared to the SOTA single-model method, GoMatching achieves substantial improvements of 39.92% and 18.77% on MOTA and IDF1, respectively.
Results on ArTVideo. We test TransDETR and GoMatching on ArTVideo to compare the zero-shot text spotting capabilities for arbitrary-shaped text. For a fair comparison, both TransDETR and GoMatching are trained on ICDAR15-video. Unlike ICDAR15-video and DSText which only have straight text, ArTVideo has a substantial number of curved text, so we report results under four settings: tracking results on both straight and curved text, spotting results on both straight and curved text, tracking results on curved text only, and spotting results on curved text only. As shown in Table 3, GoMatching outperforms TransDETR under all settings. Especially when involving an additional recognition task (end-to-end spotting) or only considering curved text, the performance advantages of GoMatching are more significant. This further confirms that the previous SOTA methods have unsatisfactory recognition capabilities and limited adaptability to complex scenarios. Furthermore, as shown in Fig. 1(b), GoMatching achieves excellent performance while significantly reducing the training budget.
@@
Table 3: Comparison between TransDETR and GoMatching on ArTVideo. ‘-Curved’ represents experimental results on curved text only. Best results are marked in bold.
Some visual results are provided in Fig. 4. It shows that GoMatching performs well on straight and curved text, and even more complex scene text.
@@
Figure 4: Visual results of video text spotting. Images from top to bottom are the results on ICDAR15-video, DSText, and ArTVideo, respectively. Text instances belonging to the same trajectory are assigned the same color.
@@
4.4 Ablation Studies
We first conduct comprehensive ablation studies on ICDAR15-video to verify the effectiveness of each component, and the experimental results are shown in Table 4. The impact of frame length on long-term association during inference is then studied, and the results are shown in Table 5.
Effectiveness of Rescoring Mechanism. To verify the effectiveness of the rescoring mechanism, we test three different scoring mechanisms: the original score from DeepSolo, the score recomputed by the rescoring head, and the fusion score from the rescoring mechanism. As shown in row 2 and row 3 of Table 4, the rescoring head can alleviate the performance degradation caused by the domain gap between ICDAR15-image and ICDAR15-video, and achieve gains of 1.25% and 0.97% on MOTA and IDF1, respectively. Moreover, as shown in row 4, we can observe that combining the knowledge of rescoring head learned from the new dataset with the prior knowledge of DeepSolo can further improve MOTA and IDF1 by 0.33% and 0.32%, respectively.
Effectiveness of LST-Matcher. In this part, we conduct three experiments to prove the effectiveness of the LST-Matcher. As shown in row 4 of Table 4, we only use LT-Matcher to associate high-score text instances in the current frame with trajectories in the tracking memory bank. In row 5, we only use ST-Matcher to associate high-score text instances in the current frame with trajectories of the previous frame. In addition, as shown in row 6, we employ both LT-Matcher and ST-Matcher to test LST-Matcher. We can easily observe that compared to LT-Matcher, LST-Matcher improves MOTA and IDF1 by 1.72% and 1.29% respectively, while compared to ST-Matcher, the improvement on MOTA and IDF1 are 1.12% and 5.1%, respectively. This proves the effectiveness of our proposed LST-Matcher, which can integrate the advantages of LT-Matcher and ST-Matcher.
Impact of the Frame Number in LT-Matcher. In Table 5, we further study and analyze the impact of the number of frames in LT-Matcher during inference. For the text spotting task, since a single frame may have a large number of text instances, sometimes reaching hundreds, excessive historical frame information would weaken the discrimination of text instance features, resulting in erroneous matching results. Therefore, we conduct a hyper-parameter search and find that the optimal frame number for long-term association in LT-Matcher is 6.
5 Conclusion
In this paper, we propose a simple yet strong baseline, termed GoMatching for video text spotting. GoMatching harnesses the talent of an off-the-shelf query-based image text spotter and only needs to tune a lightweight tracker, effectively addressing the limitations of previous SOTA methods in recognition. Specifically, we design the rescoring mechanism and LST-Matcher to adapt the image text spotter to unseen video datasets while empowering GoMatching with excellent tracking capability. Moreover, we establish a novel test set ArTVideo for the evaluation of video text spotting models on arbitrary-shaped text, filling the gap in this area. Experiments on public benchmarks and ArtVideo demonstrate the superiority of our GoMatching over previous SOTA methods in terms of both spotting accuracy and training cost.

### v2

@@ 4.1 Datasets and Evaluation Metrics
ICDAR15-video [13] is a word-level video text reading benchmark annotated with quadrilateral bounding boxes, comprising a training set of 25 videos and a test set of 24 videos. It focuses on wild scenarios, such as driving on the road, exploring shopping streets, walking in a supermarket, etc.
BOVText [10] is a large-scale, bilingual, and open-world benchmark for video text spotting, encompassing English and Chinese. The dataset is meticulously collected from YouTube and KuaiShou with different scenarios. The text box annotations are represented as quadrilaterals at the textline level.
DSText [19] is a newly proposed dataset, and focuses on dense and small text reading challenges in the video. This dataset provides 50 training videos and 50 test videos. Compared with the previous datasets, DSText mainly includes the following three new challenges: dense video texts, high-proportioned small texts, and various new scenarios, e.g., ‘Game’, ‘Sports’, etc. Similar to ICDAR15-video, DSText adopts word-level annotations, which are labeled with quadrilaterals.
ArTVideo is a novel word-level test set established in this work to evaluate the performance of arbitrary-shaped video text, which is absent in the VTS community. It contains 20 videos with about 30% curved text instances. Straight text is annotated with quadrilaterals, while curved text is annotated with polygons. More details are provided in the Appendix A.
Evaluation Metrics. To evaluate performance, we adopt three evaluation metrics commonly used in ICDAR15-video competition and DSText competition, including MOTA [31], MOTP, and IDF1 [32].
@@ 4.2 Implementation Details
In all experiments, we only use a single NVIDIA GeForce RTX 3090 (24G) GPU to train and evaluate GoMatching. As for the image text spotter in GoMatching, we apply the officially released DeepSolo [5]. During fine-tuning GoMatching on downstream video datasets, we only update the rescoring head and LST-Matcher, while keeping DeepSolo frozen. More inference settings can be seen in the Appendix B.
Training Setting. The text spotting part of GoMatching is initialized with off-the-shelf DeepSolo weights and kept frozen in all experiments. We optimize other modules on video datasets. We follow EfficientDet [33] to adopt the scale-and-crop augmentation strategy with a resolution of 1280. The batch size $T$ is 6. All frames in a batch are from the same video. Text instances with fusion scores higher than 0.3 are fed into the LST-Matcher during training. AdamW [34] is used as the optimizer. We adopt the warmup cosine annealing learning rate strategy with the initial learning rate being set to 0.00005. The loss weights $\lambda_{res}$ and $\lambda_{asso}$ are set to 1.0 and 0.5, respectively. For focal loss, $\alpha$ is 0.25 and $\gamma$ is 2.0 as in [14, 5]. The model is trained for 30k iterations on all downstream video datasets.
@@ 4.3 Comparison with State-of-the-art Methods
Results on ICDAR15-video. To evaluate the effectiveness of GoMatching on oriented video text, we conduct a comparison with the state-of-the-art methods on ICDAR15-video presented in Table 3(a). As can be seen, GoMatching ranks first in all metrics on the ICDAR15-video leaderboard. By effectively combining a robust image text spotter with a strong tracker, GoMatching improves the best performance by 5.08% MOTA, 0.75% MOTP, and 3.16% IDF1, respectively. Furthermore, owing to the substantial enhancement in recognition and tracking capabilities (details can be found in Sec. 4.4 and Appendix G), GoMatching outperforms the current SOTA single-model method TransDETR by 11.08% MOTA, 3.92% MOTP, and 7.31% IDF1, respectively.
Results on BOVText. Except for the English word recognition scenario, GoMatching can readily adapt to other video text recognition scenarios, such as Chinese text line recognition. For BOVText, which focuses on English and Chinese textline recognition, we employ the DeepSolo trained on bilingual textline datasets and then fine-tune GoMatching on BOVText. The results are presented in Table 3(b). It is evident that GoMatching achieves a new record on the BOVText dataset and surpasses previous methods significantly. GoMatching exhibits superior performance over the previous SOTA method CoText [11], with improvements of 41.5% on MOTA, 6.9% on MOTP, and 14.3% on IDF1. Such exceptional performance of GoMatching on BOVText suggests its proficiency in spotting both Chinese and English text in videos. Moreover, it can be easily extended to other languages by adapting the image spotter.
Results on DSText. We further conduct experiments on DSText with dense and small video text scenarios. Results are presented in Table 3(c). It is worth noting that most of the previous methods on the DSText leaderboard used an ensemble of multiple models and large public datasets to enhance their performance [19]. For example, TencentOCR integrates the detection results of DBNet [35] and Cascade MaskRCNN [36] built with multiple backbone architectures, combines the Parseq [37] text recognizer, and further improves the end-to-end tracking with ByteTrack [22]. DA adopts Mask R-CNN [38] and DBNet to detect text, then uses BotSORT [21] to replace the tracker in VideoTextSCM [39] and employs the Parseq model for recognition. As a single model with a frozen image text spotter, GoMatching also shows competitive performance compared to other ensembling methods on the leaderboard. GoMatching ranks first (22.83%) on MOTA, second (80.43%) on MOTP, and third (46.09%) on IDF1. Moreover, compared to the SOTA single-model method, GoMatching achieves substantial improvements of 45.46% and 19.66% on MOTA and IDF1, respectively.
Results on ArTVideo. We test TransDETR and GoMatching on ArTVideo to compare the zero-shot text spotting capabilities for arbitrary-shaped text. For a fair comparison, both TransDETR and GoMatching are trained on ICDAR15-video. Unlike ICDAR15-video and DSText which only have straight text, ArTVideo has a substantial number of curved text, so we report results under four settings: tracking results on both straight and curved text, spotting results on both straight and curved text, tracking results on curved text only, and spotting results on curved text only. As shown in Table 3(d), GoMatching outperforms TransDETR under all settings. Especially when involving an additional recognition task (end-to-end spotting) or only considering curved text, the performance advantages of GoMatching are more significant. This further confirms that the previous SOTA methods have unsatisfactory recognition capabilities and limited adaptability to complex scenarios. Furthermore, as shown in Fig. 1(b), GoMatching achieves excellent performance while significantly reducing the training budget.
Some visual results are provided in Fig. 4. It shows that GoMatching performs well on straight and curved text, and even more complex scene text. More visual results (including some failure cases) and analysis are provided in the Appendix G.
Table 1: Comparison results with SOTA methods on four distinct datasets. ‘†’ denotes that the results are collected from the official competition website. ‘*’: we use the officially released model for evaluation. ‘M-ME’ indicates whether multi-model ensembling is used. ‘Y’ and ‘N’ stand for yes and no. The best and second-best results are marked in bold and underlined, respectively.
@@
GoMatching (size:800) (ours) 68.51 77.52 76.59
GoMatching (size:1000) (ours) 72.04 78.53 80.11
GoMatching (size:1440) (ours) 70.52 78.25 78.70
Method MOTA ($\uparrow$) MOTP ($\uparrow$) IDF1 ($\uparrow$)
EAST + CRNN [10] -79.3 76.3 6.8
PSENet + CRNN [10] -17.0 79.2 31.3
DB + CRNN [10] -13.2 81.3 38.8
TransVTSpotter [10] -1.4 82.0 43.6
CoText [11] 11.4 80.3 48.3
GoMatching (ours) 52.9 87.2 62.6
@@
TransDETR [12]* N -22.63 79.73 26.43
GoMatching (ours) N 22.83 80.43 46.09
@@
4.4 Ablation Studies
We first conduct comprehensive ablation studies on ICDAR15-video to verify the effectiveness of each component. The experimental results are shown in Table 2. The impact of frame length on long-term association during inference is then studied, and the results are shown in Appendix C.
Effectiveness of Rescoring Mechanism. To verify the effectiveness of the rescoring mechanism, we test three different scoring mechanisms: the original score from DeepSolo, the score recomputed by the rescoring head, and the fusion score from the rescoring mechanism. As shown in row 2 and row 3 of Table 2, the rescoring head can alleviate the performance degradation caused by the domain gap between ICDAR15-image and ICDAR15-video, and achieve gains of 1.25% and 0.97% on MOTA and IDF1, respectively. Moreover, as shown in row 4, we can observe that combining the knowledge of rescoring head learned from the new dataset with the prior knowledge of DeepSolo can further improve MOTA and IDF1 by 0.33% and 0.32%, respectively. Appendix F contains more results.
Effectiveness of LST-Matcher. In this part, we conduct three experiments to prove the effectiveness of the LST-Matcher. As shown in row 4 of Table 2, we only use LT-Matcher to associate high-score text instances in the current frame with trajectories in the tracking memory bank. In row 5, we only use ST-Matcher to associate high-score text instances in the current frame with trajectories of the previous frame. In addition, as shown in row 6, we employ both LT-Matcher and ST-Matcher to test LST-Matcher. We can easily observe that compared to LT-Matcher, LST-Matcher improves MOTA and IDF1 by 1.72% and 1.29% respectively, while compared to ST-Matcher, the improvement on MOTA and IDF1 are 1.12% and 5.1%, respectively. In Fig. 5, we also demonstrate that using LST-Matcher can effectively mitigate the issue of ID switches caused by the strong appearance changes due to motion blur. These results validate that combining short-term and long-term information leads to more robust tracking outcomes, thereby enhancing the performance of video text spotting.
@@ 4.4 Ablation Studies
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Visual results of video text spotting. Images from top to bottom are the results on ICDAR15-video, BOVText, DSText, and ArTVideo, respectively. Text instances belonging to the same trajectory are assigned the same color.
@@
Figure 5: Visualization results of ST-Matcher and LST-Matcher. The first row shows the failure case that suffers from the ID switches issue when using only ST-Matcher, caused by missed detection and erroneous matching. The second row shows that LST-Matcher effectively mitigates this issue via both long and short term matching. Text instances in the same color represent the same IDs.
Table 3: Results of GoMatching under various training settings on the ICDAR15-video dataset. ‘Only Image Spotter’ and ‘Only Tracker’ refer to fine-tuning either the image spotter or tracker with another module fixed. ‘End-to-End’ denotes that training the image spotter and tracker in an end-to-end manner. ‘0.001’, ‘0.01’ and ‘0.1’ correspond to the ratios of the learning rate employed by the decoder of the image text spotter relative to the base learning rate. Due to constraints in training resources, we only optimize the parameters of the decoder component for the image text spotter.
Index Training Setting MOTA ($\uparrow$) MOTP ($\uparrow$) IDF1 ($\uparrow$)
1 ‘Only Tracker’ 72.04 78.53 80.11
2 First ‘Only Image Spotter’, Then ‘Only Tracker’ 70.82 78.09 79.64
3 ‘End-to-End’, Image Spotter’s Decoder (‘0.001’) 71.48 79.14 78.98
4 ‘End-to-End’, Image Spotter’s Decoder (‘0.01’) 70.15 78.17 77.67
5 ‘End-to-End’, Image Spotter’s Decoder (‘0.1’) 68.03 75.46 77.16
Different training strategies. To investigate the impacts of different training strategies on GoMatching, we establish three distinct settings on the ICDAR15-video dataset, as shown in Table 3. 1) We only fine-tune the tracker while keeping the image text spotter frozen (the first row of Table 3). 2) We first fine-tune the image text spotter on images extracted from ICDAR15-video and then further fine-tune the tracker with the image text spotter fixed (the second row of Table 3). 3) We jointly train the spotter’s decoder and tracker of GoMatching while trying different learning rates for the decoder (the last three rows of Table 3). As shown in the first two rows of Table 3, fine-tuning the image spotter on the downstream video dataset results in a performance decline compared to the default setting. This is due to two data-related factors: 1) minor variations in text content between frames in the same video lead to insufficient data diversity, causing the image text spotter to overfit more easily; and 2) image blurring from camera motion reduces the quality of data available for training the image text spotter. When the image text spotter and tracker are trained simultaneously (the last three rows), the model’s performance significantly decreases. Even with the decoder’s learning rate close to zero (i.e., 0.001), there is still a 1.13% drop in IDF1. As the decoder’s learning rate increases, the performance decline becomes more pronounced. This indicates that naive joint optimization of text spotting and tracking is challenging, likely due to conflicts between the two tasks. In future work, it is worth trying to establish larger, more diverse video text spotting datasets and to explore more effective multi-task optimization strategies.
5 Conclusion
In this paper, we propose a simple yet strong baseline, termed GoMatching, for video text spotting. GoMatching harnesses the talent of an off-the-shelf query-based image text spotter and only needs to tune a lightweight tracker, effectively addressing the limitations of previous SOTA methods in recognition. Specifically, we design the rescoring mechanism and LST-Matcher to adapt the image text spotter to unseen video datasets while empowering GoMatching with excellent tracking capability. Moreover, we establish a novel test set ArTVideo for the evaluation of video text spotting models on arbitrary-shaped text, filling the gap in this area. Experiments on public benchmarks and ArTVideo demonstrate the superiority of our GoMatching in terms of both spotting accuracy and training cost.
Acknowledgments and Disclosure of Funding
This work was supported in part by the National Key Research and Development Program of China under Grant 2023YFC2705700, in part by the National Natural Science Foundation of China under Grant U23B2048, 62076186 and 62225113, and in part by the Innovative Research Group Project of Hubei Province under Grant 2024AFA017. Dr Tao’s research is partially supported by NTU RSR and Start Up Grants. The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.
Appendix A More Details of ArTVideo
Due to the scarcity of curved text instances within existing video text spotting datasets, it is infeasible to evaluate the performance of video text spotting models on curved text. To fill this gap, we collected a test set named ArTVideo containing 20 video clips with a total of 884 frames, in which 18 videos were collected from YouTube and 2 videos from the BOVText test set. ArTVideo contains 6,526 text instances, including 4,632 straight text instances and 1,894 curved text instances, i.e., curved text accounts for about 30%. As shown in Fig. 6, we provide high-quality word-level annotations for both straight and curved text in two different annotated ways. The straight text is labeled with quadrilaterals, while for curved text, we follow the CTW1500 [40] and adopt a polygon with 14 points to annotate the text contour. More statistics about ArTVideo are shown in Fig. 7.
Figure 6: Visual examples from our ArTVideo. The straight and curved text are labeled with quadrilaterals and polygons, respectively. The same background color in different frames (columns) denotes the same instance.
Figure 7: Statistics of ArTVideo. (a) and (b) show the distribution of text instance numbers in each frame and the distribution of the text length of each instance, respectively. (c) shows the word cloud of text annotations in ArTVideo.
Appendix B More Details of Inference Settings
Since there is no training set in ArTVideo, we directly use the model trained on ICDAR15-video to evaluate the generalization ability of GoMatching to arbitrary-shaped text. The association score threshold is set to 0.2. For ICDAR15-video, we set the shorter size of the input image to 800, 1000, and 1440, with 800 aligned with the setting in TransDETR and 1440 aligned with the setting in DeepSolo. As for BOVText, DSText, and ArTVideo, the shorter sizes are set to 1000, 1280, and 1440, respectively. All the ablation studies are conducted on the setting of 1440.
Appendix C Impact of the Frame Number in LT-Matcher
In Table 5, we further study and analyze the impact of the number of frames for long-term association in LT-Matcher during inference. For the text spotting task, since a single frame may have a large number of text instances, sometimes reaching hundreds, excessive historical frame information would weaken the discrimination of text instance features, resulting in erroneous matching results. Therefore, we conduct a hyper-parameter search and find that the optimal frame number is 6.
@@ Appendix C Impact of the Frame Number in LT-Matcher
Table 5: Results of different score fusion strategies on ICDAR5-video. ‘Mean’, ‘Geo-mean’, and ‘Maximum’ denote the arithmetic mean, geometric mean, and the maximum score fusion strategies, respectively. The best results are highlighted in bold.
Strategy MOTA ($\uparrow$) MOTP ($\uparrow$) IDF1 ($\uparrow$)
Mean 70.46 78.38 78.29
Geo-mean 70.29 78.39 78.26
Maximum 70.52 78.25 78.70
Table 6: Results of using different image sizes on ICDAR15-video. ‘Size’ means the size of the shorter side of the input image during inference. The best results are highlighted in bold.
Method MOTA ($\uparrow$) MOTP ($\uparrow$) IDF1 ($\uparrow$) FPS ($\uparrow$)
TransDETR(Size: 800) 60.96 74.61 72.80 12.69
GoMatching(Size: 800) 68.51 77.52 76.59 14.41
GoMatching(Size: 1000) 72.04 78.53 80.11 10.60
Table 7: Comparison between TransDETR and GoMatching. ‘T-Para.’ and ‘A-Para.’ denote the number of all parameters and the trainable parameters in each model, respectively.
Method #T-Para. (M) #A-Para. (M)
TransDETR 39.35 39.58
GoMatching 32.79 75.38
Figure 8: Visual comparison of GoMatching with and without the rescoring mechanism. The values in parentheses indicate the confidence scores of the detected text. $\rightarrow$’ points to the filtered texts due to low confidence without using rescoring. $\rightarrow$’ points to the text whose confidence score has been improved by the rescoring mechanism. The rescoring mechanism increases the confidence scores of small texts and blurry texts, preventing them from being filtered out by the threshold and thereby cultivating a better tracking candidate pool.
Table 8: Video text detection performance on ICDAR2013-video  [karatzas2013icdar]. The best results are highlighted in bold.
Method Precision ($\uparrow$) Recall ($\uparrow$) F-measure ($\uparrow$)
Free 79.7 68.4 73.6
TransDETR 80.6 70.2 75.0
GoMatching w/o rescoring 92.4 65.7 76.8
GoMatching 89.5 74.8 81.5
Table 9: Comparison results of detection AP on the ICDAR13-video between with and without the rescoring mechanism.
Method AP $\text{AP}_{\text{S}}$ $\text{AP}_{\text{M}}$ $\text{AP}_{\text{L}}$
w/o rescoring 26.2 11.6 40.1 49.8
w/ rescoring 29.3 (+3.1) 15.5 (+3.9) 42.7 (+2.6) 51.9 (+2.1)
Appendix D Comparison of Different Score Fusion Strategies
To investigate the impact of the score fusion strategy in the rescoring mechanism, we further evaluate two other strategies: (1) the arithmetic mean score fusion strategy and (2) the geometric mean score fusion strategy, denoted as mean and geo-mean, respectively. The arithmetic mean score fusion strategy takes the average of the scores from the image text spotter and the rescoring head as the final score for each query, while the geo-mean score fusion strategy uses the geometric mean. These two strategies can be formulated as:
$c_{mean}=(c_{o}+c_{r})/2,$ (13)
$c_{geo-mean}=\sqrt{c_{o}*c_{r}},$ (14)
where $c_{mean}$ and $c_{geo-mean}$ denote the final score of the two strategies, respectively. $c_{o}$ and $c_{r}$ are the scores from the image text spotter and the rescoring head, respectively.
From Table 5, we can see that employing the maximum score fusion strategy achieves the best performance on MOTA and IDF1 among all the strategies. As for the other two strategies, extremely low confidence scores from the image text spotter may result in low final scores, probably leading to missed detections. Therefore, we adopt the maximum score fusion strategy in the rescoring mechanism of GoMatching by default.
Appendix E More Comparisons between TransDETR and GoMatching
In Table 7, we provide the results of using three different image sizes in GoMatching during inference on ICDAR15-video. In the first row of Table 7, the shorter side of the input image is set to 800, which is the same as the default setting of TransDETR. It is evident from Table 7 that GoMatching significantly outperforms TransDETR in all settings. Meanwhile, it also shows that GoMatching outperforms TransDETR under its default setting in terms of both inference speed and spotting accuracy. With the increase in image size (e.g., 1000), GoMatching offers better spotting performance at the cost of decreased inference speed.
Moreover, we compare the number of all parameters and the trainable parameters of GoMatching and TransDETR, as shown in Table 7. It is noteworthy that GoMatching requires a much smaller training budget than TransDETR owing to its simpler architecture design, making it a simple but strong baseline for video text spotting.
Figure 9: More visualization results of TransDETR and GoMatching. Failure cases of text detection and recognition are highlighted with ellipses and rectangles, respectively.
Appendix F More Results for the Rescoring Mechanism
In Table  9, we provide the video text detection results of GoMatching on ICDAR13-video  [karatzas2013icdar] and compare them with Free and TransDETR. As shown in the table, without the rescoring mechanism, GoMatching relies on the original results from DeepSolo, resulting in a 4.5% decrease on Recall and only a 1.8% improvement on F-measure compare to TransDETR. This is due to the domain gap between image data and video data, directly using an image spotter leads to a low confidence and consequently low Recall on video data. When encompasses the rescoring mechanism, GoMatching achieves a 9.1 improvement on Recall compared to DeepSolo and a 6.5% F-measure enhancement compared to TransDETR. These improvements highlight the effectiveness of rescoring mechanism in alleviating the domain gap and leading to a better tracking candidate pool. The impressive results of GoMatching are not merely attributed to the introduction of a robust image text spotter.
To further explore how the rescoring mechanism eases the domain gap, we calculate the AP of detection results on ICDAR13-video, as shown in the Table  9. Incorporating the rescoring mechanism effectively improves the detection performance of DeepSolo on video datasets, particularly for small text, resulting in a 3.9% improvement on $\text{AP}_{\text{S}}$. We also present more visual results to embody the potency of rescoring mechanism in Fig. 8.
Appendix G More Visualization Results
We present more visualization results of TransDETR and GoMatching in Fig. 9, including some failure cases. It can be observed that GoMatching exhibits a significant improvement in text recognition performance compared to TransDETR. It should be noted that GoMatching may also experience failures due to the image-video domain gap and extreme cases, such as very small text instances and significant motion blur. These issues can be mitigated by employing a stronger text spotter with a more representative backbone and training on larger, more diverse datasets.